OpenStudy (jonnyvonny):

May someone assist me with the question: Find the point on the graph of f(x)=x^2 that is closest to the point (2, (1/2)). Thank you :)

OpenStudy (jonnyvonny):

I got to the point which I took the derivative of the equation I got, leaving me with: 0=4(x^3+x-1), in which, synthetic division doesn't work, thus, I desire not to use a calculator to find its roots. I could have done something wrong, along the way to getting my answer, however.

OpenStudy (anonymous):

If you mean the vertex. It's 0,0 Although I am probably wrong.

OpenStudy (jonnyvonny):

@voce , I don't, thanks though :).

OpenStudy (jonnyvonny):

From my first post, I desire to know if I messed up, where, if not, what can I do from here?

OpenStudy (anonymous):

i think it should be \[d^2=(x-2)^2+(x^2-\frac{1}{2})^2\] which is a quadratic

OpenStudy (anonymous):

a point on the parabola is (x, x^2). your point is (2, 1/2) the distance between them is \[\sqrt{\left( x-2 \right)^2+\left( x^2-\frac{ 1 }{ 2 } \right)^2}\] minimize this by taking the derivative and setting equal to 0 and solve for x.

OpenStudy (jonnyvonny):

@satellite73 , indeed, that's what I got, then I distributed, raised the power (to eliminate the radical), then took the derivative.

OpenStudy (anonymous):

what @pgpilot326 said, although i would ignore the square root

OpenStudy (anonymous):

oh, and i was wrong of course, it is not a quadratic

OpenStudy (anonymous):

it is \[x^4-4 x+\frac{17}{4}\] i think

OpenStudy (jonnyvonny):

^, you removed it a step before me, but, this changes not the result, which I got was: \[x^4+2x^2-4x+4.25\] In which, upon taking the derivative:\[4x^3+4x-4\rightarrow 4(x^3+x-1)\]

OpenStudy (anonymous):

there is no \(2x^2\) in it

OpenStudy (anonymous):

\[2(x-2)+4x\left( x^2-\frac{ 1 }{ 2} \right)=0\]

OpenStudy (jonnyvonny):

@satellite73 , how not?

OpenStudy (anonymous):

you get \(x^2\) from \((x-2)^2\) and \(-x^2\) from \((x^2-\frac{1}{2})^2\)

OpenStudy (jonnyvonny):

Ohh, you took the derivative, and distributed not...yes?

OpenStudy (anonymous):

no i did not take the derivative yet just was saying that \[(x-2)^2+(x^2-\frac{1}{2})^2=x^4-4 x+\frac{17}{4}\]

OpenStudy (anonymous):

\[4x^3-2x+2x-4=4(x^3-1)\Rightarrow x = 1\] is the only real root

OpenStudy (anonymous):

so (1, 1) is the point on the parabola y = x^2 closest to the point (2. 1/2)

OpenStudy (anonymous):

then you see the derivative is what @pgpilot326 wrote \[4(x^3-1)\] but the \(x^2\) term drops out in the initial calculation, before taking the derivative (or after, since it is not there to begin with)

OpenStudy (anonymous):

in other words \(x^4+2x^2-4x+4.25\) is wrong, it should be \(x^4-4x+4.25\)

OpenStudy (jonnyvonny):

Mmm, yes, but how does is drop out? I distributed incorrectly?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it's easier if you don't multiply until after you take the derivative. just use the chain rule

OpenStudy (jonnyvonny):

Wait, I see the negative.

OpenStudy (anonymous):

\((x-2)^2=x^2-4x+4\) \((x^2-\frac{1}{2})^2=x^4-x^2+\frac{1}{4}\)

OpenStudy (anonymous):

or you could make this an exercise in multiplying polynomials

OpenStudy (jonnyvonny):

motha-fuking negative...dang. Icy, icy now. Wow, that TOTALLY MONKEY WRENCHED THE WHOLE OUTCOME. Damn. Thanks a lot, the both of you. and, lol @pgpilot326

OpenStudy (anonymous):

but i understand it's difficult to let go sometimes... got to find the errors and put them to rest and relieve the mind

OpenStudy (anonymous):

and yes... lol

OpenStudy (anonymous):

damn minus sign...

OpenStudy (jonnyvonny):

Such a minor factor created SUCH a wrong result. Thanks so much. I wish I could metal the both of you for your contribution. Thanks a lot.

OpenStudy (anonymous):

welcome to math

OpenStudy (jonnyvonny):

medal*

OpenStudy (anonymous):

whew... i thought thtt is some new thing the kids are doing... metalling

OpenStudy (anonymous):

or this is really an episode of scooby doo and those "damn mettling kids" thwarted the minus sign